OpenStudy (anonymous):

\[y''+y'-2y=x^2\] \[2A+(2Ax+B)-2(Ax^2+Bx+C)=x^2\] 2A=1 But why do the following equal zero? 2Ax+B=0 \[Ax^2+Bx+C=0\]

OpenStudy (helder_edwin):

shouldn't u have \[ \large x^2=2A+(2Ax+B)-2(Ax^2+Bx+C) \] \[\large =(-2A)x^2+(2A-2B)x+(2A+B-2C) \] then \[ \large -2A=1 \] \[ \large 2A-2B=0 \] \[ \large 2A+B-2C=0 \] after equating the coefficients of both polynomials?

OpenStudy (anonymous):

yep, but still....why would it be zero though? \[\large 2A-2B=0 \]

OpenStudy (helder_edwin):

because on (my) LHS there's no \(x\) so its coefficient is zero.

OpenStudy (anonymous):

wouldn't that count for the first one too?

OpenStudy (anonymous):

I'm really sorry I'm really trying to understand this.... \[\large =(-2A)x^2+(2A-2B)x+(2A+B-2C)\]

OpenStudy (helder_edwin):

the coefficients for \(x^2\) are: on LHS 1 and on RHS -2A they must be equal so 1=-2A. the same for \(x\) and the independent term

OpenStudy (anonymous):

RHS and LHS of what? you mean of the whole equation? \[2A+(2Ax+B)-2(Ax^2+Bx+C)=x^2\] Yeah that makes more sense

OpenStudy (helder_edwin):

yes. but u gotta re-arange (order, i mean) your LHS

OpenStudy (helder_edwin):

i mean, associate like terms.